OpenStudy (anonymous):

Suppose the market price is $5. The buyer who buys the first unit of output has a willingness-to-buy equal to $10; the buyer who buys the second unit of output has a willingness-to-buy equal to $9; and the buyer who buys the first unit of output has a willingness-to-buy equal to $8. Total consumer surplus is

jimthompson5910 (jim_thompson5910):

focus on `Suppose the market price is $5. The buyer who buys the first unit of output has a willingness-to-buy equal to $10` for now what's the consumer surplus for this one buyer?

OpenStudy (anonymous):

im thinking $5.00? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes, subtract the max willingness to buy/pay and the market price

jimthompson5910 (jim_thompson5910):

do the same for the other buyers then add up all the individual surpluses

OpenStudy (anonymous):

ok ...so the other buyer would be 9-5 = 4 do i add the 4 and 5 together? kind of confused

jimthompson5910 (jim_thompson5910):

and 8-5 = 3 as well

jimthompson5910 (jim_thompson5910):

so, 5+4+3 = 9+3 = 12 is your total consumer surplus

OpenStudy (anonymous):

ohhh thats what i missed ! ok got it now thank you

jimthompson5910 (jim_thompson5910):

no problem